INVESTIGATION OF THE LOW-n* RYDBERG STATES OF CALCIUM MONOFLUORIDE NEAR THE INTERSECTION WITH THE ^{2}Σ^{+} DISSOCIATIVE POTENTIAL

\underline{DANIEL S. BYUN}, JEFFREY J. KAY, VLADIMIR S. PETROVIĆ, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

Previous work in our laboratory has shown that a single repulsive state (assigned as ^{2}Σ^{+}) is responsible for predissociation of all ^{2}Σ^{+} Rydberg states of CaCl in the low-n* region. The difference between the dissociation limit of CaCl and its first ionization potential is 15,500 cm^{−1}while the difference is only 2,800 cm^{−1}in CaF, suggesting that lower Rydberg states in CaF will not be as extensively predissociated. However, we have recently observed significant predissociation in CaF. The double resonance spectra of the low-n* (n*=5–7) Rydberg states of CaF have been reinvestigated in an attempt to characterize the repulsive ^{2}Σ^{+} state responsible for the predissociation in CaF.